OpenStudy (anonymous):

How many solutions are there to the following system of equations? 5x + 3y = 7 2x -y = 2

OpenStudy (anonymous):

hey u r on a roll this evening :D slow it down

OpenStudy (anonymous):

haha i am i am

OpenStudy (anonymous):

lol

OpenStudy (ash2326):

substitute y from second equation in first to obtain x, then use it to find y it has one unique solution

OpenStudy (anonymous):

5x + 3y = 7....(1) 2x -y = 2........(2) (2) -y=2-2x y=2x-2...(2)' (2)' substitute to (1) 5x+3(2x-2)=7 11x=13 x=13/11 x substitute to (2) 2(13/11)-y=2 y=-22/26 1st solution

OpenStudy (anonymous):

5x + 3y = 7....(1) 2x -y = 2........(2) (2)x3 6x-3y=6...(3) (1)+(3) 11x=13 x=13/11 and substitute to (2) 2(13/11) - y =2 26/11 -y =2 y=4/11 this is the 2nd solution with eliminate. sorry i got wrong about y in the 1st solution